OpenStudy (anonymous):

A biased dice is such that, the probability of getting ‘k’ on the dice is k times the probability of getting 1. What is the probability of getting an odd number?

OpenStudy (anonymous):

3/6

OpenStudy (ikram002p):

6 sided normal dice ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

but in options there are answers 1/7,3/7..etc no answer is 6 in denominator

OpenStudy (anonymous):

idk whether options are wrong

OpenStudy (queelius):

First, a dice has six sides, and the outcomes are mutually exclusive. So, that means P[1] + P[2] + ... + P[6] = 1. We are given that the probability of P[k] = k * P[1] So, replace P[k] with k * P[1] in the first equation, and solve for P[1]. Then, you have the probabilities for the other outcomes. Then, solve P[1] + P[3] + P[5].

OpenStudy (anonymous):

3/6?

OpenStudy (anonymous):

k*p[1]+

OpenStudy (queelius):

I haven't did the math. I'm waiting for you to try to do it. I gave you the steps you would need to take to solve it.

OpenStudy (anonymous):

k*p[1]+k*p[3]+k*p[5]/6

OpenStudy (queelius):

Ok, first, we need to find out what P[1] is. We don't know what that is yet, but we can figure it out. We know that P[k] = k * P[1]. And, we know that P[1] + P[2] + ... + P[6] = 1. So, replace P[k], k != 2, with k * P[1]. P[1] + 2 * P[2] + 3 * P[3] + ... + 6 * P[6] = 1 Can you do the math for that?

OpenStudy (queelius):

Ooops, mistake.

OpenStudy (queelius):

I meant to type: P[1] + 2 * P[1] + 3 * P[1] + ... + 6 * P[1] = 1

OpenStudy (queelius):

That is just P[1] * (1 + 2 + 3 + 4 + 5 + 6) = 1, so P[1] is what?

OpenStudy (anonymous):

1

OpenStudy (queelius):

No, let P[1] = p. Then, we have the equation: p*(1+2+3+4+5+6)=1. Solve for p.

OpenStudy (queelius):

That step is just algebra. If you are being given these kind of problems, they would expect you to have some background in basic algebra.

OpenStudy (queelius):

First, sum the integers 1+2+3+4+5+6. What is that sum?

OpenStudy (anonymous):

21

OpenStudy (queelius):

So, 21p = 1

OpenStudy (anonymous):

21*p=1

OpenStudy (anonymous):

p=1/21

OpenStudy (queelius):

Yes. So, P[1] = 1/21. What is P[2]?

OpenStudy (anonymous):

2/21

OpenStudy (anonymous):

9/21

OpenStudy (queelius):

Yes, so, to find out P[odd number], what's next?

OpenStudy (anonymous):

3/7

OpenStudy (queelius):

Show me your work. How did you get 3/7?

OpenStudy (anonymous):

1+3+5/21=9/21=3/7

OpenStudy (queelius):

Good.

OpenStudy (anonymous):

oh i understood they given k times so we are using 2*p[2],3*p[3] etc right

OpenStudy (queelius):

No, P[3] = 3 * P[1].

OpenStudy (queelius):

As given in the problem. That is why we first had to find out what P[1] was, using the knowledge that the exhaustive sum of the mutually exclusive probabilities must sum to 1.

OpenStudy (anonymous):

yeah.i made a mistake

OpenStudy (queelius):

Understood. Good, I think you understand the problem.

OpenStudy (anonymous):

okay.can u tell me any reference for this kind of problem.means forum,books for probability

OpenStudy (queelius):

I think Googling something like "introduction probability" would be a good start. There are certainly many good books on this, also.

OpenStudy (queelius):

I have to leave for ~10 minutes, bbiab.

OpenStudy (anonymous):

ok.bye i will post another question